Question 52 Chapter 6

Jordan took a business trip from New York to Denver. She spent two days in travel, conducted business for nine days, and visited friends for five days. She incurred the following expenses:

Airfare$550 

Lodging 3,000 

Meals 900 

Entertainment of clients 750 

How much of these expenses can Jordan deduct? (Do not round intermediate calculations. Round your final answer to the nearest whole dollar value.)
Answer #1

airfare 100% deductable= 550

lodging (3000x9/16)= 1687.5

meals ((900x 11/16)/2)= 309.375

client entertainment (750/2)= 375


total deduction= $2922

answered by: LFull22